# arXiv:2603.01219v1[math.AT] 01 Mar 2026

### v1

@@ 4.5. Example: A simply connected non-formal 8-manifold
The failure of intrinsic formality in $\tilde{M}$ can be interpreted as follows:
  1. (1)
    Primary Obstruction: The operation $m_{3}$ corresponds to the triple Massey product. In the nilmanifold $M$, the relation $d\theta=\mu\wedge\nu$ forces a non-vanishing $m_{3}([\mu],[\nu],[\cdot])$ in the transferred structure.
  2. (2)
    Isotopy Modulo $k$: While $\tilde{M}$ and its formal counterpart have isomorphic cohomology rings (level $k=2$), they are not isotopic modulo $3$. The class $m_{3}]\in Harr^{3,-1}_{cyc,\,uni}(H^{*},H^{*})=H^{4,-1}(H)$ is non-zero, hence ${\mathcal{K}}_{3}(m,0)=[m_{3}]$ where $0=(m_{3}=0,\ldots,0)$ denotes the formal enhancement.
  3. (3)
@@
(In general, we have the following observation. Let $\tilde{M}$ be the resolution of a nilmanifold $M/G$. If the primary obstruction $m_{3}$ of the invariant Nomizu model $((\Lambda\mathfrak{g}^{*})^{G},d)$ is non-trivial, and the resolution map $\pi^{*}$ is a $C_{\infty}$-morphism that is injective on the generators of $m_{3}$, then $\tilde{M}$ is non-formal.)
@@
  • $x=[\mu\wedge\bar{\eta}+\bar{\mu}\wedge\eta]$
  • $y=[\nu\wedge\bar{\eta}+\bar{\nu}\wedge\eta]$
  • $z=[i\eta\wedge\bar{\eta}]$
where $\mu,\nu,\eta$ are the 1-forms of the complex Heisenberg group. The non-triviality of $m_{3}$ on these classes arises from the “hidden” variable $\theta$ in the relation $d\theta=\mu\wedge\nu$, which allows for the non-vanishing homotopy $H(x\wedge y)$.
Triple $(a,b,c)$ $m_{3}(a,b,c)$ Triple $(a,b,c)$ $m_{3}(a,b,c)$
$(x,x,y)$ $\mu\wedge\theta$ $(y,x,y)$ $2\nu\wedge\theta$
$(x,y,x)$ $2\mu\wedge\theta$ $(y,x,z)$ $\theta\wedge\eta$
$(x,y,y)$ $\nu\wedge\theta$ $(y,y,x)$ $\nu\wedge\theta$
$(x,y,z)$ $\theta\wedge\eta$ $(z,x,y)$ $\theta\wedge\eta$
$(y,x,x)$ $\mu\wedge\theta$ $(z,y,x)$ $\theta\wedge\eta$
Table 1. Non-vanishing $m_{3}$ operations in the invariant sector of $\tilde{M}$. computed via the Merkulov formula.
@@ 5.1. The algebraic curvature tensor of $\mathcal{A}$$^{\ast}$
Therefore, $\hat{F}$ defines a linear map $S^{2}S^{2}H^{r}\to{\mathbb{R}}$.
@@
$\phi\in C^{2,-1}(H^{*},H^{*})$ be Harrison 2-cochain. Assume $\varphi(x,y)=0$ unless $(\deg x,\deg y)=(0,0)\mod r$. By (4.3), we have
@@
Case b: $r$ odd. In this case $b_{r}$ must be even since the bilinear form $\langle-,-\rangle_{\varphi}$ defines a symplectic form on $H^{r}$. Hence $b_{r}=2$. Consequently, $\dim(H^{r}\odot H^{r})=\dim(H^{r}\wedge H^{r})=1$. Since the pairing $\langle-,-\rangle_{\varphi}$ is nondegenerate, the multiplication $H^{r}\odot H^{r}\to H^{2r}$ is nonzero. Since $\dim(H^{r}\odot H^{r})=1$, a nonzero map stemming out of $H^{r}\odot H^{r}$ is automatically injective and so $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$. By [11, Theorem 3.10] stating that if $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$, then for the small quotient algebra $Q_{\mathrm{small}}^{*}$ of $\mathcal{A}$$^{\ast}$, we have $Q_{\mathrm{small}}^{k}=H^{k}$ also for k=2r1k=2r-1 and k=n2r+1k=n-2r+1. Hence, by [11, Theorem 3.10], we conclude that $Q_{\mathrm{small}}^{*}=H^{*}$, i.e., $\mathcal{Q}^{\ast}$ is formal. Therefore, also $\mathcal{A}$$^{\ast}$ is formal.
@@ 6. An extension of Cavalcanti formality result
Step 2: Terms with three identical indices (Type 3-2). The remaining terms involve three indices of one type and two of another (e.g., three $i$’s and two $j$’s). Up to cyclic permutation, there are only two distinct configurations:
@@ 7. Conclusion and final remarks
1) We have defined a complete set of invariants for the rational homotopy type of a space $X$ of finite type by introducing a stratification of isotopy modulo $k$ ($k\in\mathbb{N}^{+}$) on the set of minimal $C_{\infty}$-algebra enhancements. These are expressed via the obstruction sets $\mathcal{K}_{k}(m,m^{\prime})$, which enjoy a generalized additivity (Theorem 4.12, Proposition 4.16). While it is known that the Harrison cohomology class $m_{3}]$ corresponds to the Bianchi–Massey tensor of Crowley–Nordström, it remains an open and compelling problem to express the pentagon Massey tensor of Nagy–Nordström [33] in terms of our higher obstruction classes.
@@
Acknowledgement
The author wishes to thank Università di Roma “La Sapienza” for hospitality and excellent working conditions during her visit in September 2025, where a part of this work was discussed. Shes thanks Pavel Hájek and Johannes Nordström for sending her the papers [9] and [33], respectively. She acknowledges the assistance from Google Gemini, OpenAI’s ChatGPT, and xAI’s Grok in searching several formulas in this paper. Computational assistance from xAI’s Grok in verifying the trace-free Ricci cochain, computational assistance from OpenAI’s Chatgpt in solving the Equation $\delta\phi=m_{3}$ for the proof of Proposition 6.6, and Google’s Gemini assistance in compiling Table 1 are gratefully acknowledged. Table 1 was computed with SageMath script, which was written with the help of the Google Gemini model, which we placed as Appendix at the end of this paper.
@@ Appendix A Computational Verification of $m_{3}$
Listing 1: SageMath script for $C_{\infty}$ transfer
# Setup the Rational Exterior Algebra
A.<mu, nu, theta, eta> = ExteriorAlgebra($Q$)
# Merkulov Homotopy H such that dH + Hd = Id - Pi
@@
if form == mu * nu: return theta
if form == nu * mu: return -theta
return A(0)
# Computing the transferred m3 operation
def m3(a, b, c):
return H(a*b)*c + a*H(b*c)
# Basis for the invariant sector
x, y, z = mu, nu, eta
print(f"m3(x,y,x)={m3(x,y,x)}")

### v2

@@ 4.5. Example: The Fernández-Muñoz 8-manifold
(4.67) $\alpha=[\mu\wedge\bar{\mu}],\quad\beta_{2}=[\nu\wedge\bar{\eta}],\quad\beta_{3}=[\bar{\nu}\wedge\eta]$
where $\{\mu,\nu,\theta\}$ are the 1-forms of the complex Heisenberg group and $\eta$ is the fiber form. The structure equations $d\theta=\mu\nu$ and $d\bar{\theta}=\bar{\mu}\bar{\nu}$ imply that while these classes have trivial cup products in the cohomology ring, their higher-order interactions are non-trivial.
Applying the Merkulov formula, the triple operation $m_{3}(\alpha,\beta_{2},\beta_{3})$ is computed as:
$m_{3}(\alpha,\beta_{2},\beta_{3})$ $=p\left(h(i(\alpha)i(\beta_{2}))\cdot i(\beta_{3})\right)$
$=p\left(h(-\mu\nu\bar{\mu}\bar{\eta})\cdot(\bar{\nu}\eta)\right)$
@@
This non-vanishing $m_{3}$ represents the primary obstruction to formality. In our stratification, this implies that the point $\tilde{M}$ in the moduli space of homotopy types lies in the stratum where ${\mathcal{K}}_{3}(m,0)\neq 0$, distinguishing it from its formal counterpart with the same Betti numbers.
@@ 5.1. The algebraic curvature tensor of $\mathcal{A}$$^{\ast}$
Therefore, $R$ defines a linear map $S^{2}S^{2}H^{r}\to{\mathbb{R}}$.
@@
$\phi\in C^{2,-1}(H^{*},H^{*})$ be Harrison 2-cochain. Assume $\varphi(x,y)=0$ unless $(\deg x,\deg y)=(0,0)\pmod{r}$. By (4.3), we have
@@
Case b: $r$ odd. In this case $b_{r}$ must be even since the bilinear form $\langle-,-\rangle_{\varphi}$ defines a symplectic form on $H^{r}$. Hence $b_{r}=2$. Consequently, $\dim(H^{r}\odot H^{r})=\dim(H^{r}\wedge H^{r})=1$. Since the pairing $\langle-,-\rangle_{\varphi}$ is nondegenerate, the multiplication $H^{r}\odot H^{r}\to H^{2r}$ is nonzero. Since $\dim(H^{r}\odot H^{r})=1$, a nonzero map stemming out of $H^{r}\odot H^{r}$ is automatically injective and so $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$. By [12, Theorem 3.10] stating that if $\ker(\cdot:H^{r}\odot H^{r}\to H^{2r})=0$, then for the small quotient algebra $Q_{\mathrm{small}}^{*}$ of $\mathcal{A}$$^{\ast}$, we have $Q_{\mathrm{small}}^{k}=H^{k}$ also for k=2r1k=2r-1 and k=n2r+1k=n-2r+1. Hence, by [12, Theorem 3.10], we conclude that $Q_{\mathrm{small}}^{*}=H^{*}$, i.e., $\mathcal{Q}^{\ast}$ is formal. Therefore, also $\mathcal{A}$$^{\ast}$ is formal.
@@ 6. An extension of Cavalcanti formality result
Step 2: Terms with three identical indices (Type 3-2). The remaining terms involve three indices of one type and two of another (e.g., three $i$’s and two $j$’s). Because $b^{r}\leq 2$, we are only dealing with two variables, up to cyclic permutation, there are only two distinct configurations:
@@ 7. Conclusion and final remarks
1) We have defined a complete set of invariants for the rational homotopy type of a space $X$ of finite type by introducing a stratification of isotopy modulo $k$ ($k\in\mathbb{N}^{+}$) on the set of minimal $C_{\infty}$-algebra enhancements. These are expressed via the obstruction sets $\mathcal{K}_{k}(m,m^{\prime})$, which enjoy a generalized additivity (Theorem 4.12, Proposition 4.16). While it is known that the Harrison cohomology class $m_{3}]$ corresponds to the Bianchi–Massey tensor of Crowley–Nordström [13, Theorem 4.4], it remains an open and compelling problem to express the pentagon Massey tensor of Nagy–Nordström [33] in terms of our higher obstruction classes.
@@ Appendix A Computational Verification of $m_{3}$
Listing 1: SageMath script for $C_{\infty}$ transfer on the FM 8-manifold
label=lst:sage_transfer
# 1. Define the Graded Commutative Algebra over Rational Field
# We define generators of degree 1
names = [’mu’, ’mubar’, ’nu’, ’nubar’, ’eta’, ’etabar’, ’theta’, ’thetabar’]
A.<mu, mubar, nu, nubar, eta, etabar, theta, thetabar> = GradedCommutativeAlgebra($Q$)
# 2. Define the Differential d
# d(theta) = mu*nu, d(thetabar) = mubar*nubar, others are closed
d = A.differential({theta: mu*nu, thetabar: mubar*nubar})
# 3. Define the Propagator (Homotopy Operator) h
# We define h such that d*h + h*d = id - pi.
# Here we define the mapping of coboundaries to their primitives.
@@
if form == 0: return A(0)
res = A(0)
# Extract monomials and coefficients
for mon, coeff in form.dict().items():
# Check for the Heisenberg coboundary mu*nu
if mu*nu in A(mon):
res += coeff * (theta * (A(mon) / (mu*nu)))
# Check for the conjugate coboundary mubar*nubar
elif mubar*nubar in A(mon):
res += coeff * (thetabar * (A(mon) / (mubar*nubar)))
return res
# 4. Define the Cohomology Classes (Degree 2) as per FN Theorem 3.2
alpha = mu * mubar
beta1 = nu * nubar
beta2 = nu * etabar
beta3 = nubar * eta
# 5. Merkulov Formula for m3
# m3(a, b, c) = p( i(a)*h(i(b)*i(c)) + h(i(a)*i(b))*i(c) )
# For the FM manifold, the projection p is the identity on non-boundaries.
def m3_transfer(a, b, c):
# term1: a * h(b * c)
# term2: h(a * b) * c
# Note: h(harmonic * harmonic) is usually 0 unless the product is a boundary.
val = a * h(b * c) + h(a * b) * c
return val
# 6. Computation and Output
triple = (alpha, beta2, beta3)
result = m3_transfer(*triple)
print(”—SageMathC-infinityTransfer(FNBasis)—”)
print(f”InputTriple:alpha,beta2,beta3”)
print(f”ResultinH^5:{result}”)
# 7. Verification of the m3 identity
# Check if d(result) is zero (it must be a cycle)
print(f”Istheresultacycle?{d(result)==0}”)